INAF queue observing, 2018 Mar 19-20

Observer: R. Carini, A. Rossi
Telescope Operator: G. Bechetti
Support Astronomer: D. Thompson
Instrument: MODS1+MODS2, LBC

Summary

At the sunset sky is mostly clear
instruments: all instruments work well, but MODS2 could not be used in MOS and imaging mode because the lower left quadrant doesn't work.
Night duration (12\xB0 el twilight) 2:20am- 12:35pm UT = 10 h 15m

Observing time: *9h* (90%)

Weather time loss: 0h (0%)

technical time loss * *1.1h (10%)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Calibrations flat MODS1+MODs2 slit 5.0'', binning 1x2 dual grating   mods1r.0001-0003,mods1b.0001-0006, mods2r.0001-0003,mods2b.0001-0006 mods1 b to be repeated
Calibrations flat MODS1+ MODS2 slit 1.0'', binning 1x2 dual grating   mods1r.0004-0009, mods1b.0007-0012, mods2r.0004-0009, mods2b.0007-0012 mods1 to be repeated
standard Feige34 MODS1+2 dual grating   mods1r.0013-0015, mods1b.0013-0015 ; mods2r.0013-0015, mods2b.0013-0015 for Grazian and Antoniucci
Grazian B MODS1+2 dualgrating, 1.2" 900x6/1h mods1r.0019-0021, mods1b.0016-0018 ; mods2r.0019-0021, mods2b.0016-0018 first image with airmass greater than 1.3 completed
Grazian J MODS1 dualgrating MOS 0.25/8h mods1r.0025, mods1b.0019 MODS2 not possible to use because of the bottom left bad quadrant, seeing not good, to be checked
standard pg1047_003 LBCb+r       for Dotto
Dotto 2018 EE LBCb+r BVg+RzrI     if possible to be repeated for color comparison
Dotto 2006 UF17 LBCb+r BVg+Rzri     if possible to be repeated for color comparison
D'Avanzo GRB160601A LBCb+r r-sloan filter 1800*2=1/1h -- completed
phot stand SA 104 LBCb+r BVgr+RzrI     for Dotto, D'Avanzo, Annibali
Dotto 2018 EZ LBCb+r BVg+RzrI     if possible to be repeated for color comparison
Annibali R23 NGC3738 LBCb+r g+r 1.0 /1.0h   completed
Annibali R40 NGC5477 LBC b+r g+r 14 m   last 3 images for each LBC have counts very high, be careful, but the first 4 images should be good
Calibration Sky Flat LBCb=B,r LBCr=r,z       Davanzo e Dotto, Annibali ( in r)
Calibration BIAS LBC        
Calibration arc MODS1 +2 dualgrating 1x2     for Palazzi, see log
Calibration slitflat 1" and 5" MODS1 +2 dualgrating 1x2     for Palazzi , see log
Calibration arc MODS1 +2 red+blue only grating     for Montalto some are saturated, see log
Calibration slitflat MODS1 +2 red+blue only grating     for Montalto some are saturated, see log

Details

Calibrations
  • flat slit 5.0,binning 1x2, dual grating mods1r.20180320.0001-0003,mods1b.20180320.0001-0006, mods2r.20180320.0001-0003,mods2b.20180320.0001-0006
  • mods1b.0001-0003 are saturated , we be repeated
  • flat slit 1.0'', binning 1x2, dual grating mods1r.0004-0009, mods1b.0007-0012, mods2r.0004-0009, mods2b.0007-0012
  • mods1b.0010-0012 are saturated, to be repeated

03:20 Slew to STD Feige34 MOS,dualgrating for Grazian and Antoniucci
  • acqBinoMODS feige34_ls.acq mods1r.20190320.0010-0012, mods2r.20190319.0010-0012
  • seeing 1" on guider
  • execBinoMODS feige34_ls.obs mods1r.0013-0015, mods1b.0013-0015, mods2r.0013-0015, mods2b.0013-0015

03:41 Slew to Grazian B
  • wrong guiodestar, change it and sed again
  • acqBinoMODS pointB_lsPA=-30.acq mods1r.20190320.0016-0018, mods2r.20190319.0016-0018
  • mods2 problem, lost 5 min
  • seeing 1" on guider
  • 04:07 execBinoMODS mods1r.0019-0021, mods1b.0016-0018, mods2r.0019-0021, mods2b.0016-0018
  • Since the first images have been taken with airmass greater than 1.3, we take another image
  • 03:54 seeing about 1.0''
  • seeing jumping to 1.5"

04:04 Slew to Grazian J , we try in binocular if works, but of course ods2 blue data will not be good for all masks
  • acqBinoMODS pointB_lsPA=-30.acq mods1r.20190320.0022-0024, mods2r.20190319.0022-0023
  • acquisition in mods2 did not work, 2 allignement stars are in bad quadrant
  • THERE IS NO FINDING CHART FOR this target! but we made it, had to edit and remove top left allignement star (double faint star?)
04:28 start execMODS --mods1
  • 08:07 execBinoMODS mods1r.0025, mods1b.0019
  • 04:40 seeing jumps to 1.5-2 ! we stop

04:48 Slew to Prandoni USS1
  • acqBinoMODS USS1_ls_newPiv.acq USS1_ls_newPiv_b.acq
  • checking pointing lost 5 min
  • 04:58 send acq again
  • 05:00 error in MODS1 again, problem during the preset
  • 05:08* we try again the acquisition., the seeing is too bad and it is impossibile the guiding. Some clouds are coming.

05:20 we slew to LBC Dotto program, 40 min loss due to TCS problem

06:14 start with photometric standard pg1047_003
  • dohybrid
  • 06:28 ob, seeing about 1.4''-1.5'', the sky is clear

06:33 slew to Dotto's program, target 2018 EE
  • dofpia
  • copointing
  • ob

06:56 slew to 2006 UF17
  • dofpia
  • copointing
  • ob

07:17 slew to D'Avanzo program target GRB160601A
  • dofpia
  • copointing
  • ob for photometric calibration on Panstars stars
  • 07:38 seeing 1.3" on r image
  • 07:41 ob for science
  • 07:45 DIMM 1.1"
  • 07:50 seeing 1.3" on r image
  • 08:07 seeing 1.4" on r image
  • 08:16 seeing 1.2" on r image

* slew to standard for Dotto Sa104
  • dofpia
  • 08:40 seeing 1.0" on r image
  • 08:45 seeing 0.8" on r image

08:55 slew to 2018 EZ
  • dofpia
  • copointing
  • 09:04 lost 5 minutes copointing blue image is flat, not clear why
  • 09:08 send OB science
  • seeing 1" on image

09:12 slew to Annibali R23
  • dofpia
  • copointing
  • seeing 0.8" on image
  • 09:28 run ob NGC3738_1.ob
  • 09:30 seeing about 1.0''
  • 09:54 dofpia
  • 10:01run NGC3738_2.ob
  • seeing about 0.8''
  • 10:29 dofpia
  • 10:39 NGC3738_3.ob
  • seeing about 1.0''
  • 10:57 seeing 1.1''-1.2'''

11:02 problem with the connection, maybe the problem is in the observatory because also eduroam is out of work

11:11 come back the connection
  • run the ob of calibration for R23

11:19 slew to R 40
  • dofpia

11:21 connection lost again
  • from the office of D. Thompson is possible connect with obs4
  • 11:35 the connection has come back also in the remote room, but now there is a problems in the home directory, we have to come back to Dave's office. we connect with obs2, but we can not open RB_Science. Moreover during the connection dofpia stuck. We connect to obs4 and try to open from there RB_Science, but it doesn't work.

11:53 the connection works well
  • run dofpia

12:05 copointing R40 (NGC5477)
  • run NGC5477_copointing.ob
  • the connection is unstable
  • 12:09 run NGC5477_1.ob
  • seeing about 1.0''on the DIMM

12:30 the panel of sky brightness and of the other parameters doesn't work,
  • the background of the images are too high from images at ~12:25, 17k(g-band) and 10k (r-band), we stopped
  • the first 4 images should be good
Calibrations

FLAT FIELDS blue: r, B; red: z ,r PA= 0 and 180

bias for LBC

MODS, PALAZZI_29 binning 1x2
  • arc mods1r.0026-0028 , mods1b.0020-0022,==mods2r.0024-0028 , mods2b.0021-0026==
  • flat 1" mods1r.0029-0034 , mods1b.0023-0028,
  • flat 5" mods1b.0029-0034

Montalto
  • arc mods1r.0035-0037, mods2b.0028-0030
  • the first arc of mods2 is saturated (lamp Hg), we halve the exposure time to 0.5, take the image mods2b.0031 but it is saturated again. we can decrease further texp.
  • flat mods1r.0038-0046 saturated QTH1+ND1.5 and VFLAT +clear
  • mods2b.0032-0037 saturated QTH1+ND1.5 and QTH1+ clear
  • mods1r.0044-0046 ok (Ne,Kr+Xe, Ar)
  • mods2b.0038-0040 ok (Hg,Kr+Xe,Ar)
  • take again the the saturated flats decreasing the time exposure
  • mods1r.0047-0049 no good
  • mods2b.0041-0043 no good
  • mods1.0051-0052 ok VFLAT +clear
  • mods2b.0044 no good. MODS2 stuck
  • time for calibration is over

-- %USERSIG{RobertaCarini - 2018-03-20}%

Comments

 
Topic revision: r20 - 26 Mar 2018, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback